OpenStudy (sugarmochi):

Please help me ! ! ! A Circle has a radius of 12cm. Find the percentage change in the area of the circle if the radius increases by 20%

OpenStudy (photon336):

okay so I believe that if you increase the radius by 20% so if we represent the original radius as r the new radius becomes 1.20r

OpenStudy (sugarmochi):

so... 12 x 1.20?

OpenStudy (photon336):

so I think we would approach it like this. we substitute 1.20r into the formula for the area. \[A = \pi*r^{2}\] \[A_{2} = \pi*(1.20r)^{2}\]

OpenStudy (photon336):

@sugarMochi \[A_{2} = 1.44*\pi*r^{2} | A_{1} = \pi*r^{2}\]

OpenStudy (photon336):

@SugarMochi what do you realize about the new area?

OpenStudy (sugarmochi):

44 more than the original area ?

OpenStudy (photon336):

44% more

OpenStudy (sugarmochi):

oHHHH i get it now !!!! Thankyou so so so much

OpenStudy (photon336):

do you see what I did?

OpenStudy (sugarmochi):

mhmmm yep yep

OpenStudy (photon336):

let's check again we can actually put in 12 cm and check our answer

OpenStudy (photon336):

\[(12cm)^{2}*\pi = 144\pi = area \] \[(12*1.20)^{2} = 207\pi = area \]

OpenStudy (photon336):

so if we multiply 144pi *(1.44) we end up getting 207pi

OpenStudy (photon336):

@SugarMochi we actually did not have to plug in any numbers we just used the fact that a 20% increase was 1.20 then we plugged it into the formula for area.

OpenStudy (sugarmochi):

yeessss i get it now. Thankyou lots and lots and lots :')))

OpenStudy (photon336):

no problem